OpenStudy (anonymous):

need the solutions √(5x-4)=6

OpenStudy (anonymous):

Squaring both sides of the equation gives \[(\sqrt{5x+6})^{2} = 6^{2}\] 5x+6=36 5x=30 x=6

OpenStudy (anonymous):

\[(\sqrt{5x-4})^2=6^2\]

OpenStudy (anonymous):

is not x=6

OpenStudy (anonymous):

oh woops, i just saw its a 4 and not a 6 <.< well the method still works (squaring both sides, sorry bout that.

OpenStudy (anonymous):

5x-4= 36 5x=40 x = 8

OpenStudy (anonymous):

I think the answer is x=6

OpenStudy (anonymous):

I was wrong

OpenStudy (anonymous):

u ringht is x=8

OpenStudy (anonymous):

Thanks Joemath

OpenStudy (anonymous):

after square 5x-4=36 5x = 36+4 x= 40/5 x = 8 My eyes not see good

OpenStudy (anonymous):

thanks California

OpenStudy (anonymous):

ty